Da Find the sum of the geometric Sequence (finite) Ž 8()-! = ©® Find the sixth term of (ax+3y)
Answer #1

10(a) 6 s . (5)? 6 8E b-c 3 ا= b 79) 6 8 E ). 3 6 = 24 s izi 3 3 3 2.4*! 1 + 1 +i 24 (tuhan tahun to a most ritus 8 = 1/3 n=610) (b) (20+34) ? Toti n=12 Trrito (sixth teem) ncg ист a obr s8+1=6 s8=5 a=2x b=34 5 T6 =12 (5(2x]? By = 121 (2x)7 35 y 5 1